HETAV PATEL

Grade 11,

A charge of Q coulomb is located at the centre of a cube If the corner of the cube is takenas the originthen the flux coming out from the faces of the cube in the direction of X axis will be

A charge of Q coulomb is located at the centre of a cube If the corner of the cube is takenas the originthen the flux coming out from the faces of the cube in the direction of X axis will be

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

use gauss's law

cube have 6 faces all face are symettrical so from each face equal flux will come out

Φ =∫E.ds  =Q/ε

 so flux from one face  = Q/6ε

 two face are in the x direction so total flux =Q/3ε

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin